OpenStudy (airyana1114):

Can someone help me solve this equation step-by-step? f(1) when f(x) = 5x + 1

OpenStudy (danjs):

if output = f(input) = 5*(input) +1 then y = f(x) = 5*x + 1 what about when the input =1 instead of x

OpenStudy (danjs):

y = f(1) = ...

OpenStudy (airyana1114):

so do I just replace every x to the right of the equal sign with 1?

OpenStudy (danjs):

yep, you got it

OpenStudy (airyana1114):

thank you so much :D

OpenStudy (danjs):

or with whatever they want you to evaluate f(@airyana1114 ) = 5*(@airyana1114 ) + 1

OpenStudy (danjs):

welcome

OpenStudy (airyana1114):

okay, i get it now, if I tag you in another one would you mind helping me out?

OpenStudy (danjs):

but it is a number here, so you can simplify that 5*1 + 1

OpenStudy (airyana1114):

which means y = 6